# Demystifying AI Platform Design for Distributed Inference of Next-Generation LLM models

### v2

@@ II-A LLM Architecture
Mixture of Experts (MoE)[45] are special class of LLMs that consist of multiple “expert” multi-layer perceptron (MLP) layers, denoted as ‘E’, out of which ‘K’ experts are selected for each input token. In contrast, dense language models can be considered as a special case of MoEs, where E = K = 1, meaning there is only one expert MLP layer, and it is used for every input token. By introducing multiple experts and selectively activating a subset of them for each token, MoEs can effectively scale model capacity while maintaining efficient computation and memory usage, making them a promising approach for building larger and more capable language models. Some popular MoE models are Switch Transformers[45], Mixtral 8x7B[46], Mixtral 8x22B[47], DBRX[48], GROK[20], GPT-4[49, 50, 17]. Table IV presents some of the state-of-the-art LLMs and associated dimensions.
@@ III GenZ : Generative LLM Analyzer
Model $D_{model}$ Layers
Heads
($H_{kv}$)
$W_{ff}$
Experts
Gemma2-2B [51] 2304 26 8(4) 4 -
LLaMA3-8B [52] 4096 32 32(8) 3.5 -
Gemma2-27B [51] 4608 46 32(16) 8 -
Mixtral-8x22B [47] 6144 56 48(8) 2.66 8(2)
LLaMA3-70B [52] 8192 80 64(8) 3.5 -
GPT3-175B [7] 12288 96 96 4 -
LLaMA3-405B [52] 16384 126 128(8) 3.25 -
GPT4-1.8T [49] 10752 120 84 4 16(2)
Dense-5T 49152 128 192(24) 4 -
MoE-10T 13824 128 108(12) 4 32(4)
TABLE IV: Different LLM architectures evaluated in this work.
@@ III-A Model Profiler
Table IV shows the parameters of various LLM models that we study in this work. These models serve as a representative set of current and future LLMs. GenZ model can model different LLM architectures, including MoE-based LLMs (Mixtral-8x7B [46] and GPT4) and Mamba-based models [58]. LLaMA2, LLaMA3, Mixtral, and GPT3 architecture are available openly. We represent GPT4 as a 1.8T parameter mixture-of-experts model, with 120 layers [49, 50, 17]. A single layer of GPT4 has 16 experts of  [^1] B parameters each, and two experts are activated for each token. We also hypothesize two future LLM models  [^2] We do not train these models for accuracy, which is not the focus of this work. We use it to study its computational behavior. to represent scaled-up Dense-GQA and MoE models.
@@ III-C Platform Characterizer
Figure 5: GenZ (a-d) LLM inference modeling against vLLM inference in prefill and decode stage for varying batch size on various platforms with different $\tau_{p}$. Line color blue, red, and green represent LLaMA2-7B, LLaMA2-13B, and OPT-175B.
@@ III-D2 GenZ Validation
Prefill & Decode validation: Using vLLM, we evaluate LLaMA2-7B, LLaMA2-13B, and OPT-175B across various platforms, using randomly generated dummy paragraphs as input, ranging from 500 to 2000 tokens, with each model generating a fixed 200 output tokens. Fig. 5(a)-(d) compares the models’ prefill TTFT and decode throughput on NVIDIA platform. The geomean error in prefill and decode predictions between real and GenZ-predicted values is 2.73% and 1.85%, respectively, across different models and platforms. The average empirically-measured efficiency factors used for different hardware configurations are V100: 0.45, A100: 0.4, 1×H100: 0.55, 2×H100: 0.64, 4×H100: 0.66, and 8×H100: 0.75.
@@
Platform interconnect validation: To verify the collectives time at the platform scale, we benchmark all-reduce communication primitive with nccl-tests [74], a communication primitive performance benchmark for NVIDIA GPUs. GenZ ’s collective communication times are validated with a platform size of 2 GPUs, 4 GPUs, and 8 GPUs. For collective communications, we observed an average link efficiency of 75% for NVLINK, which gives an effective link BW of 350 GB/s for each GPU in HGX box. We profile all model (Table IV for different input and output lengths and collect their all-reduce (AR) message sizes. The message size of each AR call is very small ($<$ 128 KB) for decode, while for prefill, per call message size is a few hundred MBs. Fig. 8 compares the real hardware latency for three different platform sizes and the corresponding latency generated by GenZ . Since the prefill and decode stages have stark differences in message size, we verify the collective time for prefill and decode independently. We found that for decode message sizes, the link latency, $T_{Link}$ is the dominating, thus the latency seems almost constant. While prefill AR messages are large enough, that the link bandwidth, $BW_{Link}$, is the main contributor to collective time. Collective for all datapoints, there is a 3.89% and 2.7% geomean error for decode and prefill message sizes between real values and GenZ values.
@@ IV Impact of Inference Optimizations on Hardware
LLM inference is one of the most active fields of research in recent years. This has led to a rapid rise in the number of innovations and optimizations being introduced. Table V summarizes a few of the most popular techniques and their impact on compute and memory requirements from the AI inference platform. These fall into broadly 3 buckets, i.e., 1 Model architecture change, 2 System and algorithmic optimizations which don’t change the model quality, and finally 3 Algorithmic optimization with model quality changes.
GenZ supports almost all of the techniques shown in Table V, with being the only framework that supports modeling MoEs, mamba, and hybrid models, at the same time supporting sliding window attention, speculative decoding, weight sparsity, KV pruning, and mixed precision computation and storage.
@@ 3 Lossy Model optimization with impact on model quality
Quanitization Reduced bitwidths /
@@
TABLE V: Various techniques for optimizing LLMs and their impact on compute and memory requirements.
@@ VII-A Case Study I: Scaling HW Characteristics
We study how improving different HW characteristics impacts the latency of different stages. For all our experiments, we assume a hypothetical Dense-GQA-5T parameter model defined in Table IV. We explore four key hardware characteristics: TFLOPS, Memory BW , ICN link bandwidth, ICN link latency. To understand the effect of each characteristic on LLM inference performance, we vary each parameter in an isolated fashion. We vary the context length $\in\{1k,32k\}$, and for running with chunking; we vary the chunk size $\in\{384,1536\}$ and batch size $\in\{1,256\}$. Table VI summarizes this case study’s findings and Fig. 20 quantifies the observed trends.
@@ VII-A4 ICN Link Latency Scaling
TABLE VI: Improvement in LLM Inference at different stages as we scale different platform characteristics.
@@
TABLE VII: Platform architecture designs for comparison. Each of our platform’s architectures is inspired by a current system architecture. Power of 4 x DGX B200 platform(8xB200 GPU) [82]. Assumed power of a larger ASIC platform (NPUs have 10x flops of B200 GPU). Power of 16 GroqRack[85]. The architectures in the table are mapped to GenZ parameters defined in Section III-B, Section III-C. Column 3 is $FLOPS$, col 4 has BWmem, Capmem, BWsram, Capsram, col 5 has scaling dimension along with BWlink for each network level.
@@ VII-B Case Study II: Comparing Diverse Platforms
Using GenZ , we compare four representative platform architectures defined in Table VII running various workloads for current models (8B, 70B, 405B) and future model architectures (5T, 10T). Since the size and architecture of these platforms are completely different, there are different costs associated with each system. As we can’t estimate the cost of all components, we use energy consumed as a proxy and Tokens/kWH as the comparison metric. We calculate the energy used in running workloads on different platforms. The energy consumption of each platform is modeled as a linear function of the utilization of its individual components  [98]. We consider four main power components: Static/Idle, Compute, Memory, and Network. For each operator, the total energy is:
@@ VII-C Case Study III: Exploring HBD Design Choices
The concept of a high-bandwidth domain (HBD), defined as a group of NPUs interconnected via high-bandwidth scale-up links, plays a pivotal role here. Nvidia has progressively expanded the size of HBDs, scaling from 8 NPUs in the DGX H100 system to an impressive 72 NPUs in the GB200 NVL72. This case study delves into determining the optimal HBD size and designing effective interconnections between HBDs. To this end, we explore and compare various network architectures across a 256-NPU setup. We use the same individual NPU with 9 PFlops compute and 256 GB HBM, providing 13.5 TB/s for all configurations. Table VIII shows the bandwidth and link latency of different interconnect types that we consider for building the platforms. We keep the network topology fixed as 2 levels of switch followed by a ring for the third dimension. Using these, we build five different network architectures shown in Table IX. We ran models with TP=64 and PP=4.
@@
TABLE VIII: Link latency and bandwidth for different interconnects. High bandwidth(SL) is a scale-up link similar to NVlink, UAlink, etc.  [100, 101, 102, 90, 103, 104, 105, 106, 107].
@@
TABLE IX: Comparison of different network configurations
@@ VII-E Extreme Scale Requirements: AI assistant
We envision an enormous 10T parameter Super-LLM model as the futuristic model from Table IV. Our aim is to employ the Super-LLM as a standard AI assistant capable of engaging in real-time conversations with humans, furnishing answers on diverse topics, and retaining recollections of past interactions. Hence, supporting large context windows is imperative for AI assistants. This workload can be characterized as $S_{b}=4,\tau_{p}=\text{Variable},\tau_{d}=2000$. Given its role as a personal assistant, we would operate it with a batch size of 1. For seamless conversation, we assume it can generate output at the rate at which humans read or listen, which amounts to 300 words per minute[110].
@@ X Conclusions
We also provide linear energy estimation and aggregated batching support; future works can extend these capabilities with advanced modeling (e.g., additional scheduling modules to simulate disaggregated/heterogeneous serving).

### v3

@@ II-A LLM Architecture
Mixture of Experts (MoE)[45] are special class of LLMs that consist of multiple “expert” multi-layer perceptron (MLP) layers, denoted as ‘E’, out of which ‘K’ experts are selected for each input token. In contrast, dense language models can be considered as a special case of MoEs, where E = K = 1, meaning there is only one expert MLP layer, and it is used for every input token. By introducing multiple experts and selectively activating a subset of them for each token, MoEs can effectively scale model capacity while maintaining efficient computation and memory usage, making them a promising approach for building larger and more capable language models. Some popular MoE models are Switch Transformers[45], Mixtral 8x7B[46], Mixtral 8x22B[47], DBRX[48], GROK[20], GPT-4[49, 50, 17]. Table II presents some of the state-of-the-art LLMs and associated dimensions.
@@ III-A Model Profiler
Table II shows the parameters of various LLM models that we study in this work. These models serve as a representative set of current and future LLMs. GenZ model can model different LLM architectures, including MoE-based LLMs (Mixtral-8x7B [46] and GPT4) and Mamba-based models [58]. LLaMA2, LLaMA3, Mixtral, and GPT3 architecture are available openly. We represent GPT4 as a 1.8T parameter mixture-of-experts model, with 120 layers [49, 50, 17]. A single layer of GPT4 has 16 experts of  [^1] B parameters each, and two experts are activated for each token. We also hypothesize two future LLM models  [^2] We do not train these models for accuracy, which is not the focus of this work. We use it to study its computational behavior. to represent scaled-up Dense-GQA and MoE models.
@@ III-C Platform Characterizer
Figure 5: GenZ LLM inference modeling against vLLM inference in prefill and decode stage for varying batch size on various platforms with different $\tau_{p}$. Line color blue, red, and green represent LLaMA2-7B, LLaMA2-13B, and OPT-175B.
@@ III-D2 GenZ Validation
Prefill & Decode validation: Using vLLM, we evaluate LLaMA2-7B, LLaMA2-13B, and OPT-175B across various platforms, using randomly generated dummy paragraphs as input, ranging from 500 to 2000 tokens, with each model generating a fixed 200 output tokens. Fig. 5 compares the models’ prefill TTFT and decode throughput on NVIDIA platform. The geomean error in prefill and decode predictions between real and GenZ-predicted values is 2.73% and 1.85%, respectively, across different models and platforms. The average empirically-measured efficiency factors used for different hardware configurations are V100: 0.45, A100: 0.4, 1×H100: 0.55, 2×H100: 0.64, 4×H100: 0.66, and 8×H100: 0.75.
@@
Platform interconnect validation: To verify the collectives time at the platform scale, we benchmark all-reduce communication primitive with nccl-tests [74], a communication primitive performance benchmark for NVIDIA GPUs. GenZ ’s collective communication times are validated with a platform size of 2 GPUs, 4 GPUs, and 8 GPUs. For collective communications, we observed an average link efficiency of 75% for NVLINK, which gives an effective link BW of 350 GB/s for each GPU in HGX box. We profile all model (Table II for different input and output lengths and collect their all-reduce (AR) message sizes. The message size of each AR call is very small ($<$ 128 KB) for decode, while for prefill, per call message size is a few hundred MBs. Fig. 8 compares the real hardware latency for three different platform sizes and the corresponding latency generated by GenZ . Since the prefill and decode stages have stark differences in message size, we verify the collective time for prefill and decode independently. We found that for decode message sizes, the link latency, $T_{Link}$ is the dominating, thus the latency seems almost constant. While prefill AR messages are large enough, that the link bandwidth, $BW_{Link}$, is the main contributor to collective time. Collective for all datapoints, there is a 3.89% and 2.7% geomean error for decode and prefill message sizes between real values and GenZ values.
@@ IV Impact of Inference Optimizations on Hardware
LLM inference is one of the most active fields of research in recent years. This has led to a rapid rise in the number of innovations and optimizations being introduced. Table IV summarizes a few of the most popular techniques and their impact on compute and memory requirements from the AI inference platform. These fall into broadly 3 buckets, i.e., 1 Model architecture change, 2 System and algorithmic optimizations which don’t change the model quality, and finally 3 Algorithmic optimization with model quality changes.
GenZ supports almost all of the techniques shown in Table IV, with being the only framework that supports modeling MoEs, mamba, and hybrid models, at the same time supporting sliding window attention, speculative decoding, weight sparsity, KV pruning, and mixed precision computation and storage.
@@ 3 Lossy Model optimization with impact on model quality
Quantization Reduced bit widths /
@@
TABLE IV: Various techniques for optimizing LLMs and their impact on compute and memory requirements.
@@ VII-A Case Study I: Scaling HW Characteristics
We study how improving different HW characteristics impacts the latency of different stages. For all our experiments, we assume a hypothetical Dense-GQA-5T parameter model defined in Table II. We explore four key hardware characteristics: TFLOPS, Memory BW , ICN link bandwidth, ICN link latency. To understand the effect of each characteristic on LLM inference performance, we vary each parameter in an isolated fashion. We vary the context length $\in\{1k,32k\}$, and for running with chunking; we vary the chunk size $\in\{384,1536\}$ and batch size $\in\{1,256\}$. Table V summarizes this case study’s findings and Fig. 20 quantifies the observed trends.
@@ VII-A4 ICN Link Latency Scaling
TABLE V: Improvement in LLM Inference at different stages as we scale different platform characteristics.
@@
TABLE VI: Platform architecture designs for comparison. Each of our platform’s architectures is inspired by a current system architecture. Power of 4 x DGX B200 platform(8xB200 GPU) [82]. Assumed power of a larger ASIC platform (NPUs have 10x flops of B200 GPU). Power of 16 GroqRack[85]. The architectures in the table are mapped to GenZ parameters defined in Section III-B, Section III-C. Column 3 is $FLOPS$, col 4 has BWmem, Capmem, BWsram, Capsram, col 5 has scaling dimension along with BWlink for each network level.
@@ VII-B Case Study II: Comparing Diverse Platforms
Using GenZ , we compare four representative platform architectures defined in Table VI running various workloads for current models (8B, 70B, 405B) and future model architectures (5T, 10T). Since the size and architecture of these platforms are completely different, there are different costs associated with each system. As we can’t estimate the cost of all components, we use energy consumed as a proxy and Tokens/kWH as the comparison metric. We calculate the energy used in running workloads on different platforms. The energy consumption of each platform is modeled as a linear function of the utilization of its individual components  [98]. We consider four main power components: Static/Idle, Compute, Memory, and Network. For each operator, the total energy is:
@@ VII-C Case Study III: Exploring HBD Design Choices
The concept of a high-bandwidth domain (HBD), defined as a group of NPUs interconnected via high-bandwidth scale-up links, plays a pivotal role here. Nvidia has progressively expanded the size of HBDs, scaling from 8 NPUs in the DGX H100 system to an impressive 72 NPUs in the GB200 NVL72. This case study delves into determining the optimal HBD size and designing effective interconnections between HBDs. To this end, we explore and compare various network architectures across a 256-NPU setup. We use the same individual NPU with 9 PFlops compute and 256 GB HBM, providing 13.5 TB/s for all configurations. Table VII shows the bandwidth and link latency of different interconnect types that we consider for building the platforms. We keep the network topology fixed as 2 levels of switch followed by a ring for the third dimension. Using these, we build five different network architectures shown in Table VIII. We ran models with TP=64 and PP=4.
@@
TABLE VII: Link latency and bandwidth for different interconnects. High bandwidth(SL) is a scale-up link similar to NVlink, UAlink, etc.  [100, 101, 102, 90, 103, 104, 105, 106, 107].
@@
TABLE VIII: Comparison of different network configurations
@@ VII-E Extreme Scale Requirements: AI assistant
We envision an enormous 10T parameter Super-LLM model as the futuristic model from Table II. Our aim is to employ the Super-LLM as a standard AI assistant capable of engaging in real-time conversations with humans, furnishing answers on diverse topics, and retaining recollections of past interactions. Hence, supporting large context windows is imperative for AI assistants. This workload can be characterized as $S_{b}=4,\tau_{p}=\text{Variable},\tau_{d}=2000$. Given its role as a personal assistant, we would operate it with a batch size of 1. For seamless conversation, we assume it can generate output at the rate at which humans read or listen, which amounts to 300 words per minute[110].